# On the constrained feedback linearization control based on the MILP representation of a ReLU-ANN

### v2

@@ Abstract
Index Terms:
neural networks, feedback linearization, mixed-integer, constraint satisfaction, model predictive control.
I Introduction
@@
On the other hand, with the ability to handle constraints, Model Predictive Control (MPC) constitutes a promising candidate for navigation in the new coordinates. However, although the prediction model is linear, the constraints enforced are not trivial to solve since, over a prediction horizon, they are defined implicitly by the decision variables (the predicted states and inputs), creating a computationally intensive program online. To relax this setting and achieve real-time operability, approximation-based solutions have been proposed, creating control laws with the complexity of quadratic programs (QP). The most common technique computes the actual constraint via state feedback and extends it constantly to the whole prediction horizon [11, 12]. In a less conservative manner, predictions from the previous time step are employed to approximate the exact actual constraints [13]. While both of these solutions demonstrate computational efficiency, they pose challenges in terms of stability analysis development [13]. This difficulty persists even under the assumption of online feasibility, as the proof of constraint satisfaction is limited to the applied input rather than the entire predicted state or input trajectory. Recently, probabilistic approaches have been explored to learn the constraints via Gaussian processes [14, 15], leveraging the affine structure inherent in single-input flat systems to formulate the control as a second-order cone program.
@@
In our pursuit of constructing a robust approximation of the constraints and rendering the online routine tractable, this study delves into the application of the Rectified Linear Unit (ReLU) Artificial Neural Network (ANN). The focus is on characterizing the intricate constraints induced by FL. We opt for this neural network structure due to its versatile nature as a universal approximator and, most importantly, its representation in the form of Mixed-Integer (MI) linear constraints [16, 17]. This configuration (outlined in Fig. 1) offers the advantage in that it can seamlessly integrate with other linear constraints, including those derived from Control Lyapunov Function (CLF) or Control Barrier Function (CBF) principles [18]. Moreover, within the MPC framework, the constraints can be rigorously imposed along the entire prediction horizon without resorting to additional approximations. Notably, the MI constraints retain linearity, enabling efficient resolution through Mixed-Integer Program (MIP) solvers [19]. To the best of the authors’ knowledge, this application represents a novel contribution to the literature, offering a theoretically comprehensive solution to the challenge of intricate constraints associated with FL. Different from the approaches in related works [14, 15], our constraint reformulation, while specifically tackling uncertainty-free problems, exhibits versatility in its applicability to multi-input systems. This adaptability is achievable through an approximation that maintains a bounded error via the neural network.
The remainder of the paper is structured as follows. Section II presents the constrained problem and recalls the architecture of the ReLU-ANN employed. Section III provides the procedure for reformulating the feedforward structure of the ReLU-ANN to MI linear constraints. Therein, the key ingredients for the reformulation (as listed in Fig. 1) will be addressed and followed by the integration in common optimization-based control settings. Simulation studies are provided in Section IV. Finally, Section V concludes and addresses future directions.
@@
II Preliminaries and problem statement
@@
Fig. 2 illustrates the role of $\Phi(\boldsymbol{z},\boldsymbol{v})$ in the red block, creating a linear system to control in the new coordinates. This variable change, in general, complicates the constraints (2) into a new set, restricting not only the new input $\boldsymbol{v}$ but also the state $\boldsymbol{z}$. From (2) and (4), define such constraints as:
@@
Assumption: With such an objective, we further assume that the stabilization of (3) implies that of (1). Dealing with unstable left-over dynamics after the linearization will not be addressed in this paper.
@@
Denote $n_{k}$ number of nodes in the $k$th layer, $k\in\mathcal{I}_{1,K}$ which is characterized by the weight $\boldsymbol{W}^{k}\in\mathbb{R}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n_{k}\times n_{k-1}}}$ and the bias $\boldsymbol{b}^{k}\in\mathbb{R}^{n_{k}}$. Then, the input-output relationship of the network can be calculated as follows:
$\begin{cases}\boldsymbol{y}^{k}&=\sigma(\boldsymbol{W}^{k}\boldsymbol{y}^{k-1}+\boldsymbol{b}^{k}),\,k\in\mathcal{I}_{1,K-1},\\ \boldsymbol{y}^{K}&=\boldsymbol{W}^{K}\boldsymbol{y}^{K-1}+\boldsymbol{b}^{K},\end{cases}$ (7)
@@
III Constraint characterization with MIP
@@
In this part, we show that the constraint $(\boldsymbol{z},\boldsymbol{v})\in\mathcal{V}$ as in (5) can be guaranteed by ensuring a set of MI linear constraints. Then, their formulation (the green block in Fig. 1) will be addressed. The representation’s application in control synthesis will be discussed subsequently.
III-A Constraint characterization
@@
$\;\;\;\;\;\;\;\;\;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y^{\star}=\overline{y}^{\star}}$ (12)
s.t $\begin{cases}\boldsymbol{w}^{\top}\boldsymbol{y}^{in}+\boldsymbol{b}=\overline{y}^{\star}-\underline{y}^{\star},\overline{y}^{\star}\geq 0,\underline{y}^{\star}\geq 0,&\\ \alpha\in\{0,1\},\overline{y}^{\star}\leq U\alpha,\underline{y}^{\star}\leq-L(1-\alpha),&\\ \end{cases}$
@@
where, for the ReLU-ANN $\Phi_{NN}(\boldsymbol{z},\boldsymbol{v})$, $\boldsymbol{y}^{k}\triangleq\overline{\boldsymbol{y}}^{k}$, $\boldsymbol{W}^{k},\boldsymbol{b}^{k}$ denote the output, weight matrix, and bias of its $k$th layer. $L^{k}_{j},U^{k}_{j}$ are the lower and upper bound of the $j$th node in the $k$th layer. The transformation from (11) to (12) is also referred to in the literature as the “big-M” technique. Then, the program (13d) is a Mixed-Integer Linear Program (MILP) with the real variables $\boldsymbol{y}^{k},\overline{\boldsymbol{y}}^{k},\underline{\boldsymbol{y}}^{k}$ and the binary variable $\alpha_{j}^{k}$. An illustration of the formulation (13d) is given in Fig. 3. ∎
@@
As noted in Fig. 1 and Fig. 3, the three key ingredients of the formulation (13d) include the network parameters ($\boldsymbol{W}^{k},\boldsymbol{b}^{k}$), the nodes’ bounds ($L_{j}^{k},U_{j}^{k}$) and the approximation error bound ($\boldsymbol{\epsilon}$ as in (8)). These components will be the subject of the subsequent discussion.
@@
2) The nodes’ bounds $L_{j}^{k},U_{j}^{k}$ as in (13c): As well-known in the MIP community, the “big-M” values $L_{j}^{k}$ and $U_{j}^{k}$ directly affect the solver efficiency if chosen too large, and their optimal selection, in general, still remains a topic of interest. Yet, with the focus on ReLU-ANN, several algorithms have been proposed to tighten these bounds with different levels of complexity. For a detailed survey and comparison, we send the readers to the work [16] and the references therein. In the next section, the feasibility-based bound tightening (FBBT) procedure is chosen due to its simplicity [16] and solver-free nature.
Input: The input layer’s bound $L_{i}^{0},U_{i}^{0},i\in\mathcal{I}_{1,n_{z}+m}$.
Output: $L_{j}^{k},U_{j}^{k}$, $k\in\mathcal{I}_{1,K-1}$, $j\in\mathcal{I}_{1,n_{k}}$
1 for $j=1$ to $n_{1}$  do
2       $U_{j}^{1}=b_{j}^{1}+\sum_{i=1}^{n_{k}-1}\max\{W_{ji}^{1}U_{i}^{0},W_{ji}^{1}L_{i}^{0}\}$;
3       $L_{j}^{1}=b_{j}^{1}+\sum_{i=1}^{n_{k}-1}\min\{W_{ji}^{1}U_{i}^{0},W_{ji}^{1}L_{i}^{0}\}$;
4
5 end for
6for $k=2$ to $K-1$ do
7       for $j=1$ to $n_{k}$  do
8             $U_{j}^{k}=\sum_{i=1}^{n_{k-1}}\big{(}\max\big{\{}W_{ji}^{k}\max(0,U^{k}_{i}),$
9            $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;W_{ji}^{k}\max(0,L^{k}_{i})\big{\}}\big{)}+b_{j}^{k}$;
10             $L_{j}^{k}=\sum_{i=1}^{n_{k-1}}\big{(}\min\big{\{}W_{ji}^{k}\max(0,U^{k}_{i}),$
11            $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;W_{ji}^{k}\max(0,L^{k}_{i})\big{\}}\big{)}+b_{j}^{k}$;
12
13       end for
14
15 end for
 ///// /$W_{ji}^{k}$ is the $j$th row, $i$th column entry of $\boldsymbol{W}^{k}$.
Algorithm 1 Calculate $L_{j}^{k},U_{j}^{k}$ in (13d) with FBBT.
@@ 10             $L_{j}^{k}=\sum_{i=1}^{n_{k-1}}\big{(}\min\big{\{}W_{ji}^{k}\max(0,U^{k}_{i}),$
To elaborate, we will incorporate the constraint (9) [^1] Hereinafter, when constraint (9) is implemented, we mean the implemetation of its MIP representation (13d). into two effective constraint-handling frameworks: CLF-CBF and MPC. This showcases the adaptability of the proposed method in dealing with both stability and safety constraints, along with constraints intricately linked to model predictions.
III-B Control implementation
@@
Remark 2
In the setting (15)–(17), the benefit of the FL scheme is prominent, especially from the offline design view point. This is because, with the obtained integrator dynamics (3) of which the poles have non-positive real parts, the construction of $V(\boldsymbol{z})$ is more apparent. The simple choice can be a quadratic CLF with the weight from the linear quadratic regulator. $\square$
@@
for some functions $\Phi_{a}(\boldsymbol{z}),\Phi_{b}(\boldsymbol{z})$. This form renders the setup in (15b) unnecessary, since, given a measured state feedback $\boldsymbol{z}$, such MI constraints can be replaced by a linear constraint of variable $\boldsymbol{v}$:
@@
where $\boldsymbol{A}_{d},\boldsymbol{B}_{d}$ denote the discretized model of (3). In this formulation, the constraint (20a) is no longer linear with respect to the control sequence $\boldsymbol{v}(t|kt_{s})$ due to the dependence (20b) involved in the multiplication $\Phi_{a}(\boldsymbol{z}(t|kt_{s}))\boldsymbol{v}(t|kt_{s})$. Thus, even with the affine form (18), the constraints (20) are not trivial to handle [13]. Meanwhile, with the proposed setting, subsequently, we show that the constraints remain as MI linear constraints, even when enforced over a prediction horizon as in (20).
@@
where $\boldsymbol{A}_{d},\boldsymbol{B}_{d}$ denote the discretized model of (3) with sampling time $t_{s}$, $\mathcal{X}_{z}$ is some polytopic state constraints over $\boldsymbol{z}$ and $\ell(\boldsymbol{z},\boldsymbol{v})$ is a chosen cost function. While the constraints (21c)-(21d) are linear, from Proposition 1, the MI constraint (21b) is also linear for the variables $\boldsymbol{v}(t|kt_{s}),\boldsymbol{z}(t|kt_{s})$. Thus, with a standard quadratic cost $\ell(\boldsymbol{z},\boldsymbol{v})$, one can formulate a MIQP-based MPC with (21).
To showcase the effectiveness of the methods, we will examine numerical examples in the following section.
IV Simulation study
For simplicity, hereinafter, for all ReLU-ANN approximations with $K-1$ hidden layers, each layer will have equally $n_{k}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{n}}$ nodes. The training data will be sampled from a hyperbox noted as $\mathcal{T}$ with $n_{s}$ samples in each axis. The training is conducted with the method fitrnet from [21] with default hyperparameters. Numerical specifications for the characterization are given in Table I. The training time (TT) is also reported therein. All the MIPs were solved with the Yalmip toolbox [24] and CPLEX solver [19]. The simulation studies are presented next, followed by the discussion.
IV-A Nonlinear mass-spring damper system (MSD)
@@
where $x_{1},x_{2}$ represent the displacement and the velocity of the system, respectively, and $u$ is the constrained input signal with $|u|\leq{u}_{max}=5$. The damping is described by the Stribeck friction model $s(x_{2})=(0.8+0.2e^{-100|x_{2}|})\tanh(10x_{2})+x_{2}$. The model can be written in the form of (3) with $\boldsymbol{z}=[z_{1},z_{2}]^{\top}=[x_{1},x_{2}]^{\top}$, $\boldsymbol{A}=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$, $\boldsymbol{B}=[0,1]^{\top}$, $\Phi(\boldsymbol{z},v)=v+s(z_{2})+z_{1}$ and the input constraint now becomes $|\Phi(\boldsymbol{z},v)|\leq{u}_{max}$ which will be approximated and converted to MI constraints as in Section III-A, with training points sampled from $\mathcal{T}=\{|\boldsymbol{z}|\leq 5,|v|\leq 10\}$.
TABLE I: Numerical specifications
System $K$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{n}}$ $\boldsymbol{\epsilon}$ in (8) $\kappa$ $\beta$ $n_{s}$ TT(s)
@@
IV-B Horizontal 1D quadrotor (Quad-1D)

### v3

@@ Abstract
{IEEEkeywords}
constraint satisfaction, feedback linearization, mixed-integer, neural networks.
1 Introduction
[…]
@@ 1 Introduction
On the other hand, with the ability to handle constraints, Model Predictive Control (MPC) constitutes a promising candidate for navigation in the new coordinates. However, although the prediction model is linear, the constraints enforced are not trivial to solve since, over a prediction horizon, they are defined implicitly by the decision variables (the predicted states and inputs), creating a computationally intensive program online. To relax this setting and achieve real-time operability, approximation-based solutions have been proposed, creating control laws with the complexity of quadratic programs (QP). The most common technique computes the actual constraint via state feedback and extends it constantly to the whole prediction horizon [11, 12]. In a less conservative manner, predictions from the previous time step are employed to approximate the exact actual constraints [13]. While both of these solutions demonstrate computational efficiency, they pose challenges in terms of stability analysis development [13]. This issue persists even under the assumption of online feasibility, as the proof of constraint satisfaction is limited to the applied input rather than the entire predicted state or input trajectory. Recently, probabilistic approaches have been used to learn the constraints via Gaussian processes [14, 15], leveraging the affine structure in single-input flat systems to formulate a controller as a second-order cone program.
In our pursuit of constructing a robust approximation of the constraints and rendering the online routine tractable, this study explores the application of the Rectified Linear Unit (ReLU) Artificial Neural Network (ANN). The focus is on characterizing the intricate constraints induced by FL. We opt for this neural network structure due to its versatile nature as a universal approximator and, most importantly, its reprensentability in the form of Mixed-Integer (MI) linear constraints [16, 17]. This configuration offers the advantage in that it can seamlessly integrate with other linear constraints, including those derived from Control Lyapunov Function (CLF) or Control Barrier Function (CBF) principles [18]. Moreover, within the MPC framework, the constraints can be rigorously imposed along the entire prediction horizon without resorting to additional approximations. Notably, the MI constraints retain linearity, enabling efficient resolution through Mixed-Integer Program (MIP) solvers [19]. To the best of the authors’ knowledge, this application is novel to the literature, offering a theoretically comprehensive solution to the problem of intricate constraints associated with FL. Different from the approaches in related works [14, 15], our constraint reformulation, while specifically tackling uncertainty-free problems, exhibits versatility in its applicability to multi-input systems. This adaptability is achievable through an approximation that maintains a bounded error via the ANN.
The remainder of the paper is structured as follows. Section 2 presents the constrained problem and recalls the architecture of the ReLU-ANN employed. Section 3 provides the procedure for reformulating the feedforward structure of the ReLU-ANN to MI linear constraints. Therein, the key ingredients for the reformulation (outlined in Fig. 1) will be addressed and followed by the integration in common optimization-based control settings. Simulation studies are provided in Section 4. Finally, Section 5 concludes and addresses future directions.
@@
2 Preliminaries and problem statement
@@
Fig. 2 illustrates the role of $\Phi(\boldsymbol{z},\boldsymbol{v})$ in the purple block, creating a linear system to control in the new coordinates. This variable change, in general, complicates the constraints (2) into a new set, restricting not only the new input $\boldsymbol{v}$ but also the state $\boldsymbol{z}$. From (2) and (4), define such constraints as:
@@
Assumption: With such an objective, we further assume that the stabilization of (3) implies that of (1). Dealing with unstable zero dynamics after the linearization will not be addressed in this brief paper.
@@
Denote $n_{k}$ number of nodes in the $k$th layer, $k\in\mathcal{I}_{1,K}$ which is characterized by the weight $\boldsymbol{W}^{k}\in\mathbb{R}^{{n_{k}\times n_{k-1}}}$ and the bias $\boldsymbol{b}^{k}\in\mathbb{R}^{n_{k}}$. Then, the input-output relationship of the network can be calculated as follows, for $k\in\mathcal{I}_{1,K-1}$:
$\boldsymbol{y}^{k}=\sigma(\boldsymbol{W}^{k}\boldsymbol{y}^{k-1}+\boldsymbol{b}^{k}),\boldsymbol{y}^{K}=\boldsymbol{W}^{K}\boldsymbol{y}^{K-1}+\boldsymbol{b}^{K},$ (7)
@@
3 Constraint characterization with MIP
@@
In this part, we show that the constraint $(\boldsymbol{z},\boldsymbol{v})\in\mathcal{V}$ as in (5) can be guaranteed by ensuring a set of MI linear constraints. Then, their formulation will be addressed. The representation’s application in control synthesis will be discussed subsequently.
3.1 Constraint characterization
@@
$\begin{cases}\boldsymbol{w}^{\top}\boldsymbol{y}^{in}+\boldsymbol{b}=\overline{y}^{\star}-\underline{y}^{\star},\overline{y}^{\star}\geq 0,\underline{y}^{\star}\geq 0,&\\ \alpha\in\{0,1\},\overline{y}^{\star}\leq U\alpha,\underline{y}^{\star}\leq L(\alpha-1),{y^{\star}=\overline{y}^{\star}},&\\ \end{cases}$ (12)
@@
where, for the ReLU-ANN $\Phi_{NN}(\boldsymbol{z},\boldsymbol{v})$, $\boldsymbol{y}^{k}\triangleq\overline{\boldsymbol{y}}^{k}$, $\boldsymbol{W}^{k},\boldsymbol{b}^{k}$ denote the output, weight matrix, and bias of its $k$th layer. $L^{k}_{j},U^{k}_{j}$ are the lower and upper bound of the $j$th node in the $k$th layer. The transformation from (11) to (12) is also referred to in the literature as the “big-M” technique. Then, the program (13d) is a Mixed-Integer Linear Program (MILP) with the real variables $\boldsymbol{y}^{k},\overline{\boldsymbol{y}}^{k},\underline{\boldsymbol{y}}^{k}$ and the binary variable $\alpha_{j}^{k}$. An illustration of the formulation (13d) is given in Fig. 3.
@@
As noted in Fig. 3, the three key ingredients of the formulation (13d) include the network parameters ($\boldsymbol{W}^{k},\boldsymbol{b}^{k}$), the nodes’ bounds ($L_{j}^{k},U_{j}^{k}$) and the approximation error bound ($\boldsymbol{\epsilon}$ as in (8)). These components will be discussed subsequently.
@@
2) The nodes’ bounds $L_{j}^{k},U_{j}^{k}$ as in (13c): As well-known in the MIP community, the “big-M” values $L_{j}^{k}$ and $U_{j}^{k}$ directly affect the solver efficiency if chosen too large, and their optimal selection, in general, still remains a topic of interest. Yet, with the focus on ReLU-ANN, several algorithms have been proposed to tighten these bounds with different levels of complexity. For a detailed survey and comparison, we send the readers to the work [16] and the references therein. In the next section, the feasibility-based bound tightening (FBBT) procedure is chosen due to its simplicity [16] and solver-free nature. With the input layer $\boldsymbol{y}^{0}\triangleq[\boldsymbol{z},\boldsymbol{v}]^{\top}\in\mathbb{R}^{n_{z}+m}$, the calculation will be adopted from Section 4.3 in [16].
@@
To elaborate, we will incorporate the constraint (9) [^1] Hereinafter, when constraint (9) is implemented, we mean the implementation of its MIP representation (13d). into two effective constraint-handling frameworks: CLF-CBF and MPC. This showcases the adaptability of the proposed method in dealing with both stability and safety constraints, along with constraints intricately linked to model predictions.
3.2 Control implementation
@@
Remark 1.
In the setting (15)–(17), the benefit of the FL scheme is prominent, especially from the offline design view point. This is because, with the obtained integrator dynamics (3) of which the poles have non-positive real parts, the construction of $V(\boldsymbol{z})$ is more apparent. The simple choice can be a quadratic CLF from the linear quadratic regulator. $\square$
@@
for some functions $\Phi_{a}(\boldsymbol{z}),\Phi_{b}(\boldsymbol{z})$. This form renders the setup in (15b) unnecessary, since, given a state feedback $\boldsymbol{z}$, such MI constraints can be replaced by a linear constraint of $\boldsymbol{v}$:
@@
where $\boldsymbol{A}_{d},\boldsymbol{B}_{d}$ denote the discretized model of (3). In this formulation, the constraint (20a) is no longer linear with respect to the control sequence $\boldsymbol{v}(t|kt_{s})$ due to the dependence (20b) involved in the multiplication $\Phi_{a}(\boldsymbol{z}(t|kt_{s}))\boldsymbol{v}(t|kt_{s})$. Thus, even with the affine form (18), the constraints (20) are not trivial to handle [13]. Meanwhile, with the proposed setting, we will show that the MI constraints remain linear, even when enforced over a prediction horizon as in (20).
@@
where $\boldsymbol{A}_{d},\boldsymbol{B}_{d}$ denote the discretized model of (3) with sampling time $t_{s}$, $\mathcal{X}_{z}$ is the polytopic state constraint, and $\ell(\boldsymbol{z},\boldsymbol{v})$ is a chosen cost function. While the constraints (21c)-(21d) are linear, from Proposition 1, the MI constraint (21b) is also linear for the variables $\boldsymbol{v}(t|kt_{s}),\boldsymbol{z}(t|kt_{s})$. Thus, with a standard quadratic cost $\ell(\boldsymbol{z},\boldsymbol{v})$, one can formulate a MIQP-based MPC with (21). To show the efficacy of the methods, we will examine numerical examples in the next section.
4 Simulation study
For simplicity, hereinafter, for all ReLU-ANN approximations with $K-1$ hidden layers, each layer will have equally $n_{k}={\bar{n}}$ nodes. The training data will be sampled from a hyperbox noted as $\mathcal{T}$ with $n_{s}$ samples in each axis. The training is conducted with the method fitrnet from [21] with default hyperparameters. Numerical specifications for the characterization are given in Table 1. The training time (TT) is also reported therein. All the MIPs were solved with the Yalmip toolbox [24] and CPLEX solver [19]. Next, the simulation studies will be presented, followed by the discussion.
4.1 Nonlinear mass-spring damper system (MSD)
@@
where $x_{1},x_{2}$ represent the displacement and the velocity of the system, respectively, and $u$ is the constrained input signal with $|u|\leq{u}_{max}=5$. The damping is described by the Stribeck friction model $s(x_{2})=(0.8+0.2e^{-100|x_{2}|})\tanh(10x_{2})+x_{2}$. The model can be written in the form of (3) with $\boldsymbol{z}=[z_{1},z_{2}]^{\top}=[x_{1},x_{2}]^{\top}$, $\boldsymbol{A}=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$, $\boldsymbol{B}=[0,1]^{\top}$, $\Phi(\boldsymbol{z},v)=v+s(z_{2})+z_{1}$ and the input constraint now becomes $|\Phi(\boldsymbol{z},v)|\leq{u}_{max}$ which will be approximated and converted to MI constraints as in Section 3.1, with training points sampled from $\mathcal{T}=\{|\boldsymbol{z}|\leq 5,|v|\leq 10\}$.
Table 1: Numerical specifications
System $K$ ${\bar{n}}$ $\boldsymbol{\epsilon}$ in (8) $\kappa$ $\beta$ $n_{s}$ TT(s)
@@
Figure 5: Trajectory tracking with the MPC (21) for (23).
4.2 Horizontal 1D quadrotor (Quad-1D)